# SSFO: Self-Supervised Faithfulness Optimization for Retrieval-Augmented Generation

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) systems require Large Language Models (LLMs) to generate responses that are faithful to the retrieved context. However, faithfulness hallucination remains a critical challenge, as existing methods often require costly supervision and post-training or significant inference burdens. To overcome these limitations, we introduce Self-Supervised Faithfulness Optimization (SSFO), the first self-supervised alignment approach for enhancing RAG faithfulness. SSFO constructs preference data pairs by contrasting the model’s outputs generated with and without the context. Leveraging Direct Preference Optimization (DPO), SSFO aligns model faithfulness without incurring labeling costs or additional inference burden. We theoretically and empirically demonstrate that SSFO leverages a benign form of likelihood displacement, transferring probability mass from parametric-based tokens to context-aligned tokens. Based on this insight, we propose a modified DPO loss function to encourage likelihood displacement. Comprehensive evaluations show that SSFO significantly outperforms existing methods, achieving state-of-the-art faithfulness on multiple context-based question-answering datasets. Notably, SSFO exhibits strong generalization, improving cross-lingual faithfulness and preserving general instruction-following capabilities. We release our code and model at the anonymous link: https://github.com/chkwy/SSFO
@@ 1 Introduction
With the widespread deployment of Retrieval Augmented Generation (RAG) RAG ; jokinen2024need , large language models (LLMs) openai2023gpt ; touvron2023llama are increasingly expected to generate responses that adhere closely to the provided context trust_align ; rag-hat ; ragtruth , especially in scenarios where their internal parametric knowledge is insufficient or outdated. As a result, faithfulness hallucination zhou2023context ; survey_taxonomy ; es2024ragas has emerged as a critical challenge for current LLMs.
@@
While a growing body of literature has emerged to address faithfulness hallucination, existing methods often suffer from two main drawbacks: reliance on costly supervision and post-training or increased inference burden. Current approaches generally fall into two categories: post-training-based methods and decoding strategy-based methods. Post-training-based methods trust_align ; rag-hat ; context-dpo ; liu2025chatqa leverage supervised fine-tuning and direct preference optimization dpo to enhance faithfulness. However, these methods often necessitate expensive human or advanced AI supervision (e.g., GPT-4). Meticulously creating thousands to tens of thousands of training data incurs significant annotation costs. On the other hand, decoding strategy-based methods gema2024decore ; shi2024trusting alleviate faithfulness hallucinations through a plug-and-play approach that can be easily adapted to newly developed LLMs. However, they typically double the inference burden by requiring parallel processing with perturbed and natural inputs, increasing computational complexity.
@@
To overcome these limitations, we propose Self-Supervised Faithfulness Optimization (SSFO), the first self-supervised LLM faithfulness alignment approach. SSFO offers: 1) A self-supervised faithfulness alignment framework with negligible post-training cost (hundreds of self-generated examples), which helps SSFO easily adapt to newly developed LLM. 2) No added inference burden, which is crucial for lightweight deployment on edge devices yu2024edge . To implement SSFO, we leverage the model being optimized itself to generate preference data pairs by contrasting its outputs generated when provided with context with those generated from the query alone (i.e., without context). DPO dpo training is then applied using preference data pairs to align the model toward enhanced faithfulness. Our results demonstrate that this simple yet effective method can outperform both post-training-based methods trust_align ; context-dpo ; liu2025chatqa and decoding strategy-based methods gema2024decore in faithfulness to the context.
@@
To understand the underlying mechanism of SSFO, we show that the improvement can be attributed to the likelihood displacement phenomenon LD . We provide both theoretical analysis and empirical results demonstrating that likelihood displacement transfers probability mass from parametric-based tokens to context-aligned tokens. Building on this insight, we propose a modified DPO loss function tailored to enhance this beneficial displacement. Overall, our contributions can be concluded as follows:
  • We introduce SSFO, the first self-supervised alignment method for LLM faithfulness alignment. SSFO leverages self-generated data during training, requiring no human annotations, strong AI models, or ground-truth labels; the training signal derives entirely from contrasting the model’s own parametric knowledge (as unpreferred examples) against retrieved knowledge (as preferred examples). We show that faithfulness alignment can be achieved through self-supervision without external supervision.
  • We provide a theoretical and empirical analysis of the underlying alignment mechanism of SSFO and introduce a modified DPO loss function to enhance the alignment process. We find that likelihood displacement can be beneficial in the RAG setting and that encouraging this displacement further boosts faithfulness alignment.
@@
  • We provide comprehensive evaluations across diverse LLMs and benchmarks. Results show that SSFO achieves state-of-the-art faithfulness and superior generalization, including robust cross-lingual contextual faithfulness. Moreover, since trained with only hundreds of self-generated examples, SSFO preserves LLM’s general instruction-following ability, avoiding catastrophic forgetting catastrophic .
@@
Figure 1: (a): Existing post-training methods rely on human annotators or strong AI models to build SFT or preference datasets, resulting in heavy labeling costs and lengthy post-training processes. (b) SSFO leverages the model itself to generate its own preference data: Given query $x$, it generates a context-grounded response $y_{c}^{\prime}$ (with external knowledge) and a parametric-only response $y_{p}$ (query only). SSFO can significantly reduce faithfulness hallucinations without any external supervision and incurs only negligible post-training costs.
@@ 2 Preliminaries
Faithfulness hallucination refers to the LLM’s response incorporating incorrect or made-up information that is not faithful to the provided context zhou2023context ; survey_taxonomy ; es2024ragas . Let $\pi_{\theta}$ represent the model with parameters $\theta$ which generates a response $y$ when provided with a query $x$ with provided context $c$. The response $y$ can be conceptualized as being composed of two components: $y_{c}$ corresponds to content derived from the external context $c$, whereas $y_{p}$ represents content originating from the model’s parametric knowledge. Specifically, response $y$ can be represented as the combination of these two components, i.e., $y=y_{c}\oplus y_{p}$. To construct a faithful response, we aim to minimize discrepancies between response $y$ and ideal context-grounded response $y_{c}$.
Directed Preference Optimization (DPO) dpo : RLHF is computationally expensivecheng2023adversarial ; yuan2023rrhf and can suffer from instabilitiessong2023reward ; go2023aligning . DPO can bypass both explicit reward estimation and performing reinforcement learning to learn the policy using a single maximum likelihood objective. The DPO loss is defined as:
@@
where $(x,y_{w},y_{l})$ represents a data sample from dataset $\mathcal{D}$ consisting of a prompt $x$, a preferred completion $y_{w}$, and a unpreferred completion $y_{l}$. $\pi_{\theta}$ is the policy model undergoing optimization, and reference model $\pi_{\text{ref}}$ is the original state of the model before optimization. The hyperparameter $\beta$ controls the difference between policy model $\pi_{\theta}$ and reference model $\pi_{\text{ref}}$.
Likelihood Displacement LD ; pal2024smaug ; tajwar2024preference : Likelihood displacement is a counterintuitive phenomenon observed during direct preference optimization, where the probabilities for the preferred response $\pi_{\theta}(y_{w}\mid x)$ and the unpreferred response $\pi_{\theta}(y_{l}\mid x)$ both decrease, while the margin between them widens. Since $y_{w}$ is typically the optimal response (e.g., human-written or from a superior model), this reduction is problematic. Recent work dposhift ; AlphaPO aims to alleviate this phenomenon.
@@ 3 Methodology
In this section, we describe the methodology of the proposed self-supervised Faithfulness Optimization (SSFO). SSFO leverages self-supervised data construction and preference alignment training to reduce faithfulness hallucination in language models. Our goal is to train models to prioritize faithfulness to the provided external context over their internal parametric knowledge. This prioritization is critical for robust Retrieval-Augmented Generation (RAG) systems.
@@ 3.1 Self-Supervised Preference Data Construction
Existing approaches trust_align ; rag-hat ; context-dpo leverage DPO to mitigate faithfulness hallucinations and rely on curated preference data, often from human annotators or powerful AI models like GPT-4. While effective, these approaches incur substantial data annotation costs and post-training overhead.
To address this challenge, we propose a self-supervised data construction method that avoids any external labeling or supervision, as shown in Fig.˜1 (b). Our key idea is to exploit the LLM’s own responses under different knowledge access conditions to construct the preference pairs. Specifically, we generate two types of outputs for preference optimization:
Construction of preferred response $y_{w}$: We provide the model $\pi_{\theta}$ with both the query $x$ and the external context knowledge $c$ (from MS MARCO MS_MARCO ) to construct preferred responses, i.e., $y_{w}\sim\pi_{\theta}(y\mid x,c)$. Given the known faithful hallucination of LLMs (i.e. blend internal and external information when generating responses) trust_align ; ragtruth ; faithbench , we denote the response as a composition of context-based response $y_{c}$ and parametric-knowledge-based response $y_{p}$, such that $y_{w}=y_{c}^{\prime}=y_{c}\oplus y_{p}$.
Construction of unpreferred response $y_{l}$: We provide the model with the query $x$ only, omitting the external context. The model then generates a response based solely on its parametric knowledge: $y_{p}\sim\pi_{\theta}(y\mid x)$. We denote this response as $y_{l}=y_{p}$, since it reflects the model’s internal knowledge, and is more susceptible to hallucinations due to the absence of grounding in retrieved information.
The preference data pairs $(y_{c}^{\prime},y_{p})$ thus establish the context-grounded response as the positive example of faithfulness, and the purely parametric response as the negative example.
@@ 3.2 Self-Supervised Faithfulness Optimization
We perform Direct Preference Optimization over generated preference dataset $(y_{c}^{\prime},y_{p})$ to achieve faithfulness alignment. Specifically, Given a language model $\pi_{\theta}$, we minimize the following loss:
@@
This objective encourages the model to increase the likelihood of context-grounded response $y_{c}^{\prime}$ while penalizing the ungrounded, parametric-only response $y_{p}$. Since $y_{p}$ also contributes to the composite response $y_{c}^{\prime}$, this optimization inherently suppresses hallucinated content in $y_{c}^{\prime}$ and amplifies the influence of the grounded component $y_{c}$ within $y_{c}^{\prime}$. Through this formulation, we steer the model to reduce faithfulness hallucination (i.e. align more closely with retrieved context) by maximizing the discrepancy between outputs generated with external knowledge and those generated without it, without requiring costly external supervision.
In practice, training on just a few hundred such pairs yields significant improvements in faithfulness, outperforming methods that rely on human or strong AI-generated training data (Section˜4.3).
3.3 Analyzing Likelihood Displacement in Self-Supervised Faithfulness Optimization
Empirical studies Section˜4.1 show that although $y_{c}^{\prime}$ is an imperfect golden answer generated by $\pi_{\text{ref}}$, training with Eq.˜2 leads to a policy model $\pi_{\theta}^{*}$ that can significantly outperform $\pi_{\text{ref}}$. We attribute this performance improvement to a benign form of likelihood displacement LD ; pal2024smaug ; tajwar2024preference . Specifically, we demonstrate that in the context-based question-answering setting (i.e. RAG setting), SSFO facilitates this benign displacement by reallocating probability mass from tokens associated with parametric knowledge to those grounded in external contextual information.
s Refer to caption
Figure 2: Left: Log-likelihood of preferred response $\pi_{\theta}(y_{c}^{\prime}|x,c)$ versus unpreferred responses $\pi_{\theta}(y_{p}|x,c)$ over the course of SSFO optimization. Right: We compare the base model and optimized model on MemoTrap memotrap dataset and show the mean change in probability for context-based tokens $\Delta P(z_{c})$, and for parametric-only tokens $\Delta P(z_{p})$, revealing that optimization increases $\Delta P(z_{c})$ while decreasing $\Delta P(z_{p})$, $r$ denotes the Pearson correlation coefficient.
As shown in Fig.˜2 (Left), we observe a likelihood displacement phenomenon during optimization: the optimized model $\pi_{\theta}^{*}\text{ satisfies }P_{\pi_{\theta}^{*}}(y_{c}^{\prime}|x,c)<P_{\pi_{\theta}}(y_{c}^{\prime}|x,c)$ and $P_{\pi_{\theta}^{*}}(y_{p}|x,c)<P_{\pi_{\theta}}(y_{p}|x,c)$, i.e. probability mass is driven away from both the composite response $y_{c}^{\prime}$ and the parametric-based response $y_{p}$.
To understand where probability mass goes and ensure analytical tractability, we analyze the instantaneous update to the next-token distribution under gradient flow. Building on Theorem 5 of LD , let $W_{z}(t)$ denote the unembedding vector of token $z$ at training time $t$. The instantaneous change in the log-probability of token $z$, conditioned on input context $(x,c)$, is given by:
@@ 3.3 Analyzing Likelihood Displacement in Self-Supervised Faithfulness Optimization
Here, $W_{\text{token}(y_{c}^{\prime})}(t)$ is the unembedding vector of the token that the model is likely to generate given the context, while $W_{\text{token}(y_{p})}(t)$ corresponds to the token likely generated from the parametric-knowledge. In other words, the larger the inner product $\langle W_{z}(t),W_{\text{token}(y_{c}^{\prime})}(t)-W_{\text{token}(y_{p})}(t)\rangle$, the more positive the change in $\pi_{\theta(t)}(z|x,c)$.
Let $V(t)=W_{\text{token}(y_{c}^{\prime})}(t)-W_{\text{token}(y_{p})}(t)$ denote the direction vector. We analyze how different types of output tokens $z$ probability varied by examining the inner product $\langle W_{z}(t),V(t)\rangle$.
  • Faithful Token $z_{c}$ (derived from context $c$): With the LLM’s inherent ability to follow external context RAG ; zhou2023context ; gao2023retrieval , when generating the preferred response $y_{c}^{\prime}$ conditioned on $c$, the model is highly likely to produce tokens consistent with the context. Thus, $W_{\text{token}(y_{c}^{\prime})}$ is expected to be well aligned with $W_{z_{c}}$. In contrast, since $y_{p}$ reflects ungrounded, parametric-only generation, $W_{z_{c}}$ is likely misaligned with $W_{\text{token}(y_{p})}$. Therefore, the inner product $\langle W_{z_{c}}(t),V(t)\rangle$ is expected to be large.
@@
Therefore, in the RAG setting, the likelihood displacement mechanism preferentially transfers probability mass towards tokens $z_{c}$ that are consistent with the external context $c$. This constitutes a benign likelihood displacement, actively promoting faithfulness by reinforcing context-aligned generation while suppressing tokens derived from parametric knowledge or irrelevant content.
@@
3.4 Empirical Validation of Benign Likelihood Displacement
Setting. Our experiments utilize the MemoTrap dataset memotrap , designed to evaluate whether language models exhibit memorization traps. MemoTrap consists of instructions prompting the model to complete well-known proverbs with endings that deviate from the common completion. For instance, given the prompt "Write a quote that ends in the word ’right’: If you want a thing done right, do it __", the instructed target completion is "right". In this context, the token "right" represents external knowledge token $z_{c}$, while the commonly memorized completion token "yourself" is considered to be based on parametric knowledge token $z_{p}$.
We observe a benign form of likelihood displacement. As shown in Fig.˜2(Right), the probability of the faithful token $z_{c}$ increases after SSFO training. Furthermore, this rise is mirrored by a complementary fall in the probability of the parametric token $z_{p}$, producing a pronounced negative Pearson correlation (r=0.58r=-0.58). Probabilities for all remaining vocabulary tokens remain essentially unchanged and show no discernible correlation with $z_{c}$. A case study illustrating this displacement is presented in Fig.˜3.
3.5 Encouraging Benign Probability Displacement with SSFO-$\lambda$
As established in the previous analysis, the SSFO framework naturally induces a benign form of likelihood displacement, wherein probability mass shifts away from responses relying on parametric knowledge ($y_{p}$) toward those grounded in external context ($y_{c}$). This mechanism enhances model faithfulness by amplifying context-aligned generation. To further promote this desirable effect, we introduce SSFO-$\lambda$, a simple yet effective variant that explicitly encourages this displacement.
In contrast to prior work pal2024smaug ; dposhift ; AlphaPO ; Cal-dpo that focuses on mitigating likelihood displacement where $\pi_{\theta}(y_{w}\mid x)$ decreases, and even finds that encouraging the likelihood displacement may cause the model to exhibit crash behavior, producing unreadable responses dposhift . However, in Section˜3.3, we show that in the context-based question answering setting, a benign form of probability displacement can be encouraged to enhance the models’ faithfulness to the provided context through self-supervised preference optimization.
As proofed in dposhift , we introduce a parameter $\lambda>1$ to encourage the likelihood displacement during optimization:

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) systems require Large Language Models (LLMs) to generate responses that are faithful to the retrieved context. However, faithfulness hallucination remains a critical challenge, as existing methods often require costly supervision and post-training, or imposing significant inference burdens. To overcome these limitations, we introduce Self-Supervised Faithfulness Optimization (SSFO), a self-supervised alignment approach for enhancing faithfulness. SSFO constructs preference data pairs by contrasting the model’s outputs generated with context versus without context. Leveraging Direct Preference Optimization (DPO), SSFO aligns model faithfulness without incurring labeling costs or additional inference burdens. We analyze this faithfulness alignment process and provide empirical evidence that it leverages a benign form of likelihood displacement, shifting probability mass from parametric-based tokens to context-aligned tokens. Based on this insight, we adapt the DPO loss using a weighting scheme that encourages likelihood displacement. Comprehensive evaluations show that SSFO significantly outperforms existing methods, achieving state-of-the-art results in faithfulness on multiple context-based question-answering datasets. Notably, SSFO exhibits strong generalization, improving cross-lingual faithfulness while preserving general instruction-following capabilities. The code is available at: https://anonymous.4open.science/r/SSFO
@@ 1 Introduction
With the widespread deployment of Retrieval Augmented Generation (RAG) Lewis et al. (2020); Jokinen (2024), Large Language Models (LLMs) Achiam et al. (2023); Touvron et al. (2023) are increasingly expected to generate responses that adhere closely to the provided context Song et al. (2025, 2024); Niu et al. (2024). However, an LLM’s parametric knowledge from pre-training can interfere with the provided context and lead the model to generate unsupported information, known as faithfulness hallucination Zhou et al. (2023); Huang et al. ; Es et al. (2024). It has emerged as a critical challenge for current LLMs, especially in scenarios where their parametric knowledge is insufficient or outdated.
@@
A growing body of work has emerged to address faithfulness hallucination. Current approaches can be broadly categorized as follows: (1) post-training-based methods Song et al. (2025, 2024); Bi et al. (2025); Liu et al. (2025) employ supervised fine-tuning and direct preference optimization Rafailov et al. (2023) to enhance faithfulness. However, these methods often necessitate costly human or stronger LLM supervision (e.g., GPT-4). Meticulously creating thousands to tens of thousands of training examples incurs significant annotation costs. (2) decoding strategy-based methods Gema et al. (2024); Shi et al. (2024) alleviate faithfulness hallucinations through a plug-and-play approach that can be easily adapted to newly developed LLMs. However, they typically double the inference computation by requiring parallel processing with perturbed and natural inputs.
@@
To overcome these limitations, we propose Self-Supervised Faithfulness Optimization (SSFO). SSFO offers two advantages: (1) a self-supervised faithfulness alignment framework with a minor post-training cost (hundreds of self-generated examples), which helps SSFO adapt easily to newly developed LLMs. (2) no additional inference burden, which is crucial for lightweight deployment on edge devices Yu et al. (2024). To generate the training signal for faithfulness, we leverage the model’s own differential behavior when its knowledge access is altered. As shown in Fig.˜1, we leverage the model itself to generate pairs of preference data: the preferred response is generated from the query with retrieved context, while the dispreferred response is generated from the query alone, relying solely on the model’s parametric knowledge. We then apply DPO Rafailov et al. (2023) training to align the model toward enhanced faithfulness. Our results show that SSFO attains contextual faithfulness comparable to both post-training-based methods Song et al. (2025); Bi et al. (2025); Liu et al. (2025) and decoding strategy-based methods Gema et al. (2024).
@@
To understand the underlying mechanism of self-supervised faithfulness alignment, we show that it can be attributed to the likelihood displacement phenomenon Razin et al. (2025). We provide both a gradient-based analysis and empirical results demonstrating that likelihood displacement transfers probability mass from parametric-based tokens to context-aligned tokens, making the alignment well-grounded. Building on this insight, we adapt the DPO loss with a weighting scheme (SSFO-$\lambda$) to enhance this beneficial displacement and strengthen faithfulness alignment.
@@
Results show that SSFO-$\lambda$ achieves state-of-the-art faithfulness, improving performance by an average of 12% on LLaMA-3 and 27% on Mistral across faithfulness metrics relative to the instruct baseline. SSFO and SSFO-$\lambda$ also deliver superior generalization, improving cross-lingual contextual faithfulness across diverse LLMs. Moreover, since trained on only hundreds of self-generated examples, they preserve LLM’s general instruction following ability and avoid the catastrophic forgetting common in more extensive fine-tuning  Kirkpatrick et al. (2017); Dong et al. (2023, 2024).
Overall, our contributions can be summarized as follows:
  • We introduce SSFO, a self-supervised method for LLM faithfulness alignment. SSFO leverages self-generated data during training, requiring no human annotations, superior LLM models, or ground-truth labels; the training signal derives entirely from contrasting the model’s own parametric knowledge (as dispreferred examples) against retrieved knowledge (as preferred examples). We show that faithfulness alignment can be achieved via self-supervision.
  • We analyze the alignment process through the lens of likelihood displacement and provide empirical evidence that probability mass shifts from parametric to context-grounded tokens. Motivated by this finding, we investigate an easy-to-implement variant (SSFO-$\lambda$) that explicitly encourages likelihood displacement and further boosts faithfulness alignment.
  • We conduct comprehensive evaluations across diverse LLMs and benchmarks. Results show that SSFO achieves state-of-the-art faithfulness and superior generalization, including robust cross-lingual contextual faithfulness. Moreover, since trained with only hundreds of self-generated examples, SSFO preserves LLM’s general instruction-following ability, avoiding the catastrophic forgetting common in more extensive fine-tuning.
@@
Figure 1: (a): Existing post-training methods rely on human annotators or superior LLM models to construct SFT or preference datasets, resulting in heavy labeling costs and lengthy post-training processes. (b) SSFO leverages the model itself to generate preference data: Given query $x$, it generates a context-grounded response $y_{c}^{\prime}$ (with external knowledge) and a parametric-based response $y_{p}$ (query only). SSFO reduces faithfulness hallucination without external supervision and incurs negligible post‐training costs.
@@ 2 Preliminaries
Direct Preference Optimization (DPO) Rafailov et al. (2023): RLHF is computationally expensive Cheng et al. (2023); Yuan et al. (2023) and can suffer from instabilities Song et al. (2023); Go et al. (2023). DPO bypasses both explicit reward estimation and performing reinforcement learning to learn the policy using a single maximum likelihood objective. The DPO loss is defined as:
@@
where $(x,y_{w},y_{l})$ represents a data sample from dataset $\mathcal{D}$ consisting of a prompt $x$, a preferred completion $y_{w}$, and an dispreferred completion $y_{l}$. $\pi_{\theta}$ is the policy model undergoing optimization, and reference model $\pi_{\text{ref}}$ is the original state of the model before optimization. The hyperparameter $\beta$ controls the difference between policy model $\pi_{\theta}$ and reference model $\pi_{\text{ref}}$.
Likelihood Displacement Razin et al. (2025); Pal et al. (2024); Tajwar et al. (2024): Likelihood displacement is a counterintuitive phenomenon observed during direct preference optimization, where the probabilities for the preferred response $\pi_{\theta}(y_{w}\mid x)$ and the dispreferred response $\pi_{\theta}(y_{l}\mid x)$ both decrease, while the margin between them widens. Since $y_{w}$ is typically the (almost) optimal response (e.g., human-written or from a superior model), this reduction is problematic. Recent work Yang et al. (2025); Gupta et al. (2025) aims to alleviate this phenomenon.
@@ 3 Methodology
In this section, we describe the methodology of the proposed Self-Supervised Faithfulness Optimization (SSFO). SSFO leverages self-supervised data construction and preference alignment training to reduce faithfulness hallucination in language models. Our goal is to train models to prioritize faithfulness to the provided external context over their internal parametric knowledge. This prioritization is critical for robust RAG systems.
@@ 3.1 Self-Supervised Preference Data Construction
Existing approaches Song et al. (2025, 2024); Bi et al. (2025) employ DPO to mitigate faithfulness hallucinations and rely on curated preference data, often from human annotators or superior LLM models like GPT-4, as shown in Fig.˜1 (a). Although effective, these approaches incur substantial data annotation costs and post-training overhead.
To address this challenge, we propose a self-supervised data construction method that avoids external labeling or supervision, as shown in Fig.˜1 (b). Our key idea is to exploit the LLM’s own responses under different knowledge-access conditions to construct preference pairs. Specifically, we generate two types of outputs for preference optimization:
Construction of preferred response: We provide the model $\pi_{\theta}$ with the query $x$ and the retrieved context $c$ to construct preferred responses, i.e., $y\sim\pi_{\theta}(\cdot\mid x,c)$. Given the known faithfulness hallucination of LLMs (i.e., blend parametric knowledge and external context when generating responses) Song et al. (2025); Niu et al. (2024); Bao et al. (2024), we denote this partially faithful response as $y_{c}^{\prime}$.
Construction of dispreferred response: We provide the model with the query $x$ only, omitting the external context $c$. The model generates a response based solely on its parametric knowledge: $y\sim\pi_{\theta}(\cdot\mid x)$. We denote this response as $y_{p}$, which reflects the model’s internal knowledge and is more susceptible to hallucinations due to the absence of grounding in retrieved information.
The preference data pairs $(y_{c}^{\prime},y_{p})$ thus establish the context-grounded response as the positive example, and the parametric knowledge-based response as the negative example.
@@ 3.2 Self-Supervised Faithfulness Optimization
We perform DPO on the generated preference dataset $(y_{c}^{\prime},y_{p})$ to achieve faithfulness alignment. Specifically, given a language model $\pi_{\theta}$, we minimize the following loss:
@@
This objective encourages the model to increase the likelihood of the context-grounded response $y_{c}^{\prime}$ while penalizing the parametric knowledge-based response $y_{p}$. The underlying principle is that $y_{c}^{\prime}$, generated when conditioned on the external context, is generally more faithful than $y_{p}$, which relies solely on the model’s internal parametric knowledge. By widening this preference margin, the model learns to prioritize contextual information over its internal knowledge, thereby mitigating faithfulness hallucinations without costly external supervision.
In practice, training on a few hundred instances yields significant improvements in faithfulness, outperforming methods that rely on human or superior LLM-generated training data (Section˜4.3).
3.3 Analyzing and Encouraging Likelihood Displacement in Self-Supervised Faithfulness Optimization
Empirical studies (Section˜4.1) show that although $y_{c}^{\prime}$ is an imperfect answer generated by $\pi_{\text{ref}}$, training with SSFO leads to a policy model $\pi_{\theta}^{*}$ that can significantly outperform $\pi_{\text{ref}}$. We attribute these gains to a benign form of likelihood displacement Razin et al. (2025); Pal et al. (2024); Tajwar et al. (2024). Specifically, we demonstrate that in the context-based question-answering setting (i.e., RAG setting), SSFO shifts probability mass from tokens associated with parametric knowledge to those grounded in external contextual information. This effect suppresses the parametric component in both $y_{c}^{\prime}$ and $y_{p}$, favoring tokens grounded in the external context.
As shown in Fig.˜2 (left), we observe a likelihood displacement phenomenon during optimization: the optimized model $\pi_{\theta}^{*}\text{ satisfies }P_{\pi_{\theta}^{*}}(y_{c}^{\prime}|x,c)<P_{\pi_{\theta}}(y_{c}^{\prime}|x,c)$ and $P_{\pi_{\theta}^{*}}(y_{p}|x,c)<P_{\pi_{\theta}}(y_{p}|x,c)$, i.e., probability mass is driven away from both the composite response $y_{c}^{\prime}$ and the parametric response $y_{p}$.
Figure 2: Left: Log-likelihood of preferred response $\pi_{\theta}(y_{c}^{\prime}|x,c)$ versus dispreferred responses $\pi_{\theta}(y_{p}|x,c)$ over the course of SSFO optimization. Right: We compare the base instruct model and optimized model on MemoTrap Liu & Liu (2023) dataset and show the mean change for context-based tokens $z_{c}$ and parametric-based tokens $z_{p}$, revealing that optimization increases $\Delta P(z_{c})$ while decreasing $\Delta P(z_{p})$, $r$ denotes the Pearson correlation coefficient.
To understand where probability mass goes and ensure analytical tractability, we analyze the instantaneous update to the next-token distribution under gradient flow. Building on Theorem 5 of Razin et al. (2025), the instantaneous change in the log-probability of an arbitrary token $z$ from vocabulary, conditioned on input context $(x,c)$, is given by:
@@ 3.3 Analyzing and Encouraging Likelihood Displacement in Self-Supervised Faithfulness Optimization
Here, $W_{z}(t)$ denote the unembedding vector of token $z$ at training time $t$, while $W_{\text{token}(y_{c}^{\prime})}(t)$ is the unembedding vector of the token that the model is likely to generate given the context and $W_{\text{token}(y_{p})}(t)$ corresponds to the token likely generated from the parametric-knowledge. In other words, the larger the inner product $\langle W_{z}(t),W_{\text{token}(y_{c}^{\prime})}(t)-W_{\text{token}(y_{p})}(t)\rangle$, the more positive the change in $\pi_{\theta(t)}(z|x,c)$.
Let $V(t)=W_{\text{token}(y_{c}^{\prime})}(t)-W_{\text{token}(y_{p})}(t)$ denote the direction vector. We analyze how the probabilities of different types of output tokens $z$ vary by examining the inner product $\langle W_{z}(t),V(t)\rangle$.
  • Faithful Token $z_{c}$ (derived from context $c$): With the LLM’s inherent ability to follow external context Lewis et al. (2020); Zhou et al. (2023); Gao et al. , when generating the preferred response $y_{c}^{\prime}$ conditioned on $c$, the model is highly likely to produce tokens consistent with the context. Thus, $W_{\text{token}(y_{c}^{\prime})}$ is expected to be well aligned with $W_{z_{c}}$. In contrast, since $y_{p}$ reflects ungrounded, parametric-based generation, $W_{z_{c}}$ is likely unaligned with $W_{\text{token}(y_{p})}$. Therefore, the inner product $\langle W_{z_{c}}(t),V(t)\rangle$ is expected to be large.
@@
Therefore, in the Eq.˜2 setting, the likelihood displacement mechanism preferentially transfers probability mass towards tokens $z_{c}$ that are consistent with the external context $c$. This constitutes a benign likelihood displacement, actively promoting faithfulness by reinforcing context-aligned generation while suppressing tokens derived from parametric knowledge or irrelevant content.
@@
3.3.1 Empirical Validation of Benign Likelihood Displacement
Setting. Our experiments utilize the MemoTrap dataset Liu & Liu (2023), designed to evaluate whether language models exhibit memorization traps. MemoTrap consists of instructions prompting the model to complete well-known proverbs with endings that deviate from the common completion. For instance, given the prompt "Write a quote that ends in the word ’right’: If you want a thing done right, do it __", the instructed target completion is "right". In this context, the token "right" represents the external knowledge token $z_{c}$, while the commonly memorized completion token "yourself" is considered to be based on the parametric knowledge token $z_{p}$.
The SSFO optimization induces a benign form of likelihood displacement. As shown in Fig.˜2 (Right), the probability of the faithful token $z_{c}$ increases after SSFO training. Furthermore, this rise is mirrored by a complementary fall in the probability of the parametric token $z_{p}$, producing a pronounced negative Pearson correlation (r=0.58r=-0.58). Probabilities for all remaining vocabulary tokens remain essentially unchanged and show no discernible correlation with $z_{c}$. A case study illustrating this displacement is presented in Fig.˜3.
3.3.2 Encouraging Benign Probability Displacement with SSFO-$\lambda$
As established in the previous analysis, the SSFO framework induces a benign form of likelihood displacement, in which probability mass shifts away from responses that rely on parametric knowledge to those grounded in the external context. To further promote this desirable effect, we introduce SSFO-$\lambda$, a variant that explicitly encourages this displacement through a single tuning parameter. The method is easy to implement, requiring only a rescaling of the DPO objective.
Prior approaches have mainly treated likelihood displacement as a drawback Pal et al. (2024); Yang et al. (2025); Gupta et al. (2025); Xiao et al. (2024), since their “preferred” response $y_{w}$ is typically a high-quality, "golden" example (e.g., human-written), where reducing likelihood would indeed be harmful. However, our work explores using an imperfect, "silver" preferred response $y_{c}^{\prime}$ generated by the reference model itself. As analyzed in Section˜3.3, in this context-based question answering setting, encouraging the likelihood displacement proves to be an advantage for enhancing the model’s faithfulness. Motivated by Yang et al. (2025), we introduce a scaling factor $\lambda>1$ to encourage the likelihood displacement during optimization: